The equation for the ”thermite” reaction between aluminum and
iron(III)oxide is
2Al(s) + Fe2O3(s) → 2Fe(l) + Al2O3(s)
What mass of aluminum powder would be needed to react completely
with 16g of dried iron (III) oxide?
1
Expert's answer
2020-06-05T05:54:07-0400

Solution:

The balanced chemical equation:

2Al(s) + Fe2O3(s) → 2Fe(l) + Al2O3(s) 

According to the chemical equation: n(Al)/2 = n(Fe2O3)


Moles of Fe2O3 = n(Fe2O3) = Mass of Fe2O3 / Molar mass of Fe2O3

The molar mass of Fe2O3 is 160 g/mol.

n(Fe2O3) = (16 g) / (160 g/mol) = 0.1 mol


n(Al) = 2 × n(Fe2O3) = 2 × 0.1 mol = 0.2 mol

Moles of Al = n(Al) = Mass of Al / Molar mass of Al

The molar mass of Al is 27 g/mol.

Mass of Al = n(Al) × M(Al) = 0.2 mol × 27 g/mol = 5.4 g

Mass of Al is 5.4 g.


Answer: 5.4 g of aluminum powder (Al) would be needed.